# Separable models for dynamic signed networks

### v1

@@ 7 Separable 2-layer temporal models
Model (6) inherits the separability property within each time interval from Model 3 in a way that $\boldsymbol{\xi}$ and $\boldsymbol{\zeta}$ are conditionally independent given $\mathbf{y}_{t-1}$ and $\mathbf{x}_{t}$ (and therefore $\mathbf{x}^{\mathcal{F}}_{t}$ and $\mathbf{x}^{\mathcal{P}}_{t}$) in each time interval. This means that while the formation and sign layers are modelled as conditionally independent within each time interval, Model (6) does not necessarily imply temporal independence across time. In fact, the structure of $\mathbf{x}_{t}$ may be influenced not only by $\mathbf{x}_{t-1}$, but also by $\mathbf{z}_{t-1}$, and similarly, $\mathbf{z}_{t}$ may depend not only on $\mathbf{z}_{t-1}$ but also on $\mathbf{x}_{t-1}$. In other words, the presence or absence of an interaction at time $t-1$ may explain not only the formation or dissolution of an edge at time $t$, but also changes in its sign; likewise, the sign structure at $t-1$ may influence not only whether sign persist or switch, but also whether new edges appear or existing ones disappear. This form of cross-layer temporal dependence is a key feature of the model, and reflects the fact that interaction and sign dynamics are often intertwined in practice. To achieve full separability over time between the two processes, we need to assume that at any time point $t,$ $\mathbf{x}^{\mathcal{F}}$ and $\mathbf{x}^{\mathcal{P}}$ (and therefore $\mathbf{x}$) do not depend on $\mathbf{z}_{t-1},$ i.e., at any time point $t,$ $\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)=\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{x}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)=\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{x}_{t-1}\right).$
@@
Figure 3 provides a graphical representation of the separable 2-layer TERGM and illustrates that, given the interaction processes $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, the corresponding formation and persistence interaction processes – $\mathbf{x}_{1}^{\mathcal{F}}$, $\mathbf{x}_{2}^{\mathcal{F}}$, $\mathbf{x}_{1}^{\mathcal{P}}$, and $\mathbf{x}_{2}^{\mathcal{P}}$ – are fixed. This implies that $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{F}}$ are conditionally independent, as are $\boldsymbol{\xi}^{\mathcal{P}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}.$ However, temporal dependence between $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\xi}^{\mathcal{P}}$, as well as between $\boldsymbol{\zeta}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}$, still holds.
@@
Figure 3: Graphical representation of separable 2-layer TERGMs for signed network data conditional on an initial network state (not included).
@@
Table 1 presents the structure of all possible dyadic transitions in the signed network process over a given time interval, capturing how the relationships between node pairs evolve in terms of sign and connectivity.
@@
Table 1: (a) Possible transitions of a single signed edge variable $Y_{ij}$ from $t-1$ to $t;$ (b) corresponding transitions according to the separable model defined in Equation (6).
@@
$Y_{i,j,t-1}$rightarrow $Y_{i,j,t}$ $Y_{i,j,t-1}$ $Y_{i,j}^{\mathcal{F}}$ $Y_{i,j}^{\mathcal{P}}$ $Y_{i,j,t}$
$X_{i,j,t-1}$ $Z_{i,j,t-1}$rightarrow $X_{i,j}^{\mathcal{F}}$ $Z_{i,j}^{\mathcal{F}}$ $X_{i,j}^{\mathcal{P}}$ $Z_{i,j}^{\mathcal{P}}$rightarrow $X_{i,j,t}$ $Z_{i,j,t}$
00 00 00 00 00 00
00 $+1$ 00 $1$ $+1$ 00 $1$ $+1$
00 $-1$ 00 $1$ $-1$ 00 $1$ $-1$
$+1$ 00 $1$ $+1$ $1$ $+1$ 00 00
$+1$ $+1$ $1$ $+1$ $1$ $+1$ $1$ $+1$ $1$ $+1$
$+1$ $-1$ $1$ $+1$ $1$ $+1$ $1$ $-1$ $1$ $-1$
$-1$ 00 $1$ $-1$ $1$ $-1$ 00 00
$-1$ $+1$ $1$ $-1$ $1$ $-1$ $1$ $+1$ $1$ $+1$
$-1$ $-1$ $1$ $-1$ $1$ $-1$ $1$ $-1$ $1$ $-1$
@@ 7.1 Model terms
As with any ERGM-based modeling framework, the selection of network statistics is highly flexible and context-dependent. Commonly used statistics, such as edges, homophily, and transitivity, are typically included in the model to capture key structural features. STERGMs are specified by two potentially different sets of network statistics for formation and persistence.
The network statistics $\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1}\right)$ include standard binary effects. These effects are based on the assumption that the layers $\mathbf{z}^{\mathcal{F}}$ and $\mathbf{z}^{\mathcal{P}}$ are endogenous, i.e., structures that are determined within the generative process itself, rather than being specified by external covariate information, while $\mathbf{x}^{\mathcal{F}}$, $\mathbf{x}^{\mathcal{P}}$, and $\mathbf{y}_{t-1}$ are treated as exogenous. The network statistics $\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)$ for the separable 2-layer TERGM (6) assume that the layers $\mathbf{x}^{\mathcal{F}}$ and $\mathbf{x}^{\mathcal{P}}$ are endogenous, conditional on $\mathbf{y}_{t-1}$.
@@
As clearly demonstrated by Fritz et al. (2022), using lagged statistics that combine past and present edges can misrepresent the effects of SBT. For example, treating the presence of mutual friends or enemies at time $t-1$ as an exogenous covariate influencing the sign structure at time $t$ overlooks the fact that such covariates may not persist at time $t$, and may evolve in conjunction with other tie signs to form new balanced configurations. For this reason, in order to effectively model SBT effects, it is necessary to use endogenous network statistics that are implicitly dynamic – that is, they capture temporal dependencies without explicitly referencing past time points. It is important to emphasise that this reasoning also applies in our context, where SBT effects are estimated conditionally on the interaction structure. Indeed, in our case, assuming that such covariates remain constant would imply that both the interactions and their associated signs persist from time $t-1$ to time $t$, and are linked by an underlying structural dependency.
To test SBT effects, we adopt the geometrically-weighted statistics proposed by Fritz et al. (2022). These statistics can be included as binary terms $\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}}\right)$ and $\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}}\right)$ in separable 2-layer TERGMs (6) as the edgewise shared partner configurations shown in Figure 4 are used conditionally on the presence of interactions, resulting in binary configurations within a constrained support. For example, the geometrically weighted distribution of negative edgewise shared friends, $\texttt{gwesf}^{-}(\mathbf{y}_{t},\alpha)$, where $\alpha$ is the decay parameter, is equivalent to the geometrically weighted non-edgewise shared partners term, $\texttt{gwnsp}(\mathbf{z}_{t};\mathbf{x}_{t},\alpha)$, measured on $\mathbf{z}_{t}$ given $\mathbf{x}_{t}.$
@@ 8 Bayesian inference
We adopt a fully probabilistic Bayesian framework to estimate the separable 2-layer TERGM by defining a prior distribution over the model parameters $\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}})$, such that the posterior distribution is given by:
@@
Algorithm 1 AEA sampling from $\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})$
  Initialise $\boldsymbol{\zeta}^{\mathcal{F}(0)},\boldsymbol{\zeta}^{\mathcal{P}(0)}.$
  Compute $\mathbf{x}_{t}^{\mathcal{P}}$ and $\mathbf{x}_{t}^{\mathcal{P}},\forall t=1,\dots,T.$
  for $i=1,\dots,I$ do
     $\mathbf{i.}$ Propose $\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\sim h(\cdot)$
     for $t=1,\dots,T$ do
        a. Simulate $\mathbf{z}^{\mathcal{F}^{\prime}}\sim p\left(\cdot\mid\mathbf{x}_{t}^{\mathcal{F}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}}\right)$
        b. Simulate $\mathbf{z}^{\mathcal{P}^{\prime}}\sim p\left(\cdot\mid\mathbf{x}_{t}^{\mathcal{P}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{P}}\right)$
        c. Set $\mathbf{z}_{t}^{\prime}=\mathbf{z}_{t}^{\mathcal{P}^{\prime}}\cup\left(\mathbf{z}_{t}^{\mathcal{F}^{\prime}}\;\backslash\;\mathbf{z}_{t}\right)$
     end for
     $\mathbf{ii.}$ Set $(\boldsymbol{\zeta}^{\mathcal{F}(i)},\boldsymbol{\zeta}^{\mathcal{P}(i)})=(\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}})$ with probability:
$\min\left(1,\prod_{t=1}^{T}\frac{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\times\frac{\pi\left(\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\pi\left(\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\right).$ (7)
  end for
To address the sampling challenges arising from highly correlated parameters—including, as anticipated, correlations between the formation and persistence parameter groups—we employed an adaptive Metropolis–Hastings (AMH) proposal scheme (Haario, Saksman and Tamminen, 2001). Although the adaptive direction sampling scheme (used by default in Caimo and Friel (2011)) necessitates multiple chains and is less efficient over the long run, it remains valuable for sampling initial parameter values during the burn-in period.
@@
A key advantage of this approach is that network sampling from the likelihood (step a and b in Algorithm 1) can be conducted with significantly fewer auxiliary iterations and lower computational cost, thanks to the sampling space constraint and reduced complexity of the statistics which are binary.

### v2

@@ 7 Layered separable temporal models
Model (5) inherits the separability property within each time interval from Model (2) in a way that, in each interval [t1,t][t-1,t] the parameters $\boldsymbol{\xi}$ and $\boldsymbol{\zeta}$ are conditionally independent given the current interaction structure $\mathbf{x}_{t}$ and the past information $\mathbf{y}_{t-1}.$ This means that while the formation and sign layers are modelled as conditionally independent within each time interval, Model (5) does not necessarily imply full temporal independence across time. In fact, the structure of $\mathbf{x}_{t}$ may be influenced not only by $\mathbf{x}_{t-1}$, but also by $\mathbf{z}_{t-1}$, and similarly, $\mathbf{z}_{t}$ may depend not only on $\mathbf{z}_{t-1}$ but also on $\mathbf{x}_{t-1}$. In other words, the presence or absence of an interaction at time t1t-1 may explain not only the formation or dissolution of an edge at time $t$, but also changes in its sign; likewise, the sign structure at t1t-1 may influence not only whether sign persist or switch, but also whether new edges appear or existing ones disappear. This form of cross-layer temporal dependence is a key feature of the model, and reflects the fact that interaction and sign dynamics are often intertwined in practice. To achieve full separability over time between the two processes, we need to assume that at any time point $t,$ $\mathbf{x}^{\mathcal{F}}$ and $\mathbf{x}^{\mathcal{P}}$ (and therefore $\mathbf{x}$) do not depend on $\mathbf{z}_{t-1},$ i.e., at any time point $t,$ $\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)=\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{x}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)=\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{x}_{t-1}\right).$
@@
Figure 3 provides a graphical representation of the 2-layer STERGM and illustrates that, given the interaction processes $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, the corresponding formation and persistence interaction processes, $\mathbf{x}_{1}^{\mathcal{F}}$, $\mathbf{x}_{2}^{\mathcal{F}}$, $\mathbf{x}_{1}^{\mathcal{P}}$, and $\mathbf{x}_{2}^{\mathcal{P}}$, are fixed. This implies that $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{F}}$ are conditionally independent, as are $\boldsymbol{\xi}^{\mathcal{P}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}.$ However, the temporal dependence between $\boldsymbol{\xi}^{\mathcal{F}}$ and $\boldsymbol{\xi}^{\mathcal{P}}$, as well as between $\boldsymbol{\zeta}^{\mathcal{F}}$ and $\boldsymbol{\zeta}^{\mathcal{P}}$, still holds.
@@
Figure 3: Graphical representation of 2-layer STERGMs for signed network data conditional on an initial network state (not included).
@@
Table 1 presents the structure of all possible dyadic transitions in the signed network process over a given time interval, capturing how the relationships between node pairs evolve in terms of sign and connectivity. We can therefore interpret the conditional probability of observing a positive relation between nodes $i$ and $j$ at time $t$ in either the formation or persistence process by examining the corresponding log-odds expression:
$\log\left(\frac{\Pr(Z^{\mathcal{U}}_{ij,t}=+1\mid\mathbf{y}^{\mathcal{U}}_{-(ij),t},X^{\mathcal{U}}_{ij,t}=1,\mathbf{y}_{t-1})}{\Pr(Z^{\mathcal{U}}_{ij,t}=-1\mid\mathbf{y}^{\mathcal{U}}_{-(ij),t},X^{\mathcal{U}}_{ij,t}=1,\mathbf{y}_{t-1})}\right)=\boldsymbol{\zeta}^{\mathcal{U}^{\top}}\Delta_{ij}\mathbf{s}(\mathbf{z}_{t}^{\mathcal{U}};\mathbf{x}_{t}^{\mathcal{U}},\mathbf{y}_{t-1}),\text{ for }\mathcal{U}=\{\mathcal{F},\mathcal{P}\},$ (6)
where $Z^{\mathcal{U}}_{ij,t}$ is the potential sign between $i$ and $j$ at time $t$ in either formation or dissolution process, given that the interaction at time $t$ exists in the corresponding process ($X^{\mathcal{U}}_{ij,t}=1$); and the rest of the signed network in the respective process ($\mathbf{y}^{\mathcal{U}}_{-(ij),t}$). The right-hand side of the Equation (6) is the network change statistics that would result from toggling the sign of $Z_{ij,t}$, weighted by the parameter $\boldsymbol{\zeta}^{\mathcal{F}}$ or $\boldsymbol{\zeta}^{\mathcal{P}}$, respectively. A higher log-odds means the formation or persistence of the sign of the interaction is more likely under the model, based on how well it matches the network configurations encouraged by the corresponding effects.
@@
Table 1: (a) Possible transitions of a single signed edge variable $Y_{ij}$ from t1t-1 to $t;$ (b) corresponding transitions according to the separable model defined in Equation (5). When no interaction is present $(X_{ij}=0)$ the associated sign $Z_{ij}$ is not identifiable, which in turn does not allow us to model possible changes in its value.
@@
$Y_{ij,t-1}$ $\rightarrow$ $Y_{ij,t}$ $Y_{ij,t-1}$ $Y_{ij}^{\mathcal{F}}$ $Y_{ij}^{\mathcal{P}}$ $Y_{ij,t}$
$X_{ij,t-1}$ $Z_{ij,t-1}$ $\rightarrow$ $X_{ij}^{\mathcal{F}}$ $Z_{ij}^{\mathcal{F}}$ $X_{ij}^{\mathcal{P}}$ $Z_{ij}^{\mathcal{P}}$ $\rightarrow$ $X_{ij,t}$ $Z_{ij,t}$
0 0 0 0 0 0
0 +1+1 0 11 +1+1 0 11 +1+1
0 1-1 0 11 1-1 0 11 1-1
+1+1 0 11 +1+1 11 +1+1 0 0
+1+1 +1+1 11 +1+1 11 +1+1 11 +1+1 11 +1+1
+1+1 1-1 11 +1+1 11 +1+1 11 1-1 11 1-1
1-1 0 11 1-1 11 1-1 0 0
1-1 +1+1 11 1-1 11 1-1 11 +1+1 11 +1+1
1-1 1-1 11 1-1 11 1-1 11 1-1 11 1-1
@@ 7.1 Model terms
As with any ERGM-based modelling framework, the selection of network statistics is highly flexible and context-dependent. Commonly used statistics, such as edges, homophily, and transitivity, are typically included in the model to capture key structural features. STERGMs are specified by two potentially different sets of network statistics for formation and persistence.
The network statistics $\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1}\right)$ include standard binary effects. These effects are based on the assumption that the layers $\mathbf{z}^{\mathcal{F}}$ and $\mathbf{z}^{\mathcal{P}}$ are endogenous, i.e., structures that are determined within the generative process itself, rather than being specified by external covariate information, while $\mathbf{x}^{\mathcal{F}}$, $\mathbf{x}^{\mathcal{P}}$, and $\mathbf{y}_{t-1}$ are treated as exogenous. The network statistics $\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)$ and $\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)$ for the 2-layer STERGM (5) assume that the layers $\mathbf{x}^{\mathcal{F}}$ and $\mathbf{x}^{\mathcal{P}}$ are endogenous, conditional on $\mathbf{y}_{t-1}$.
@@
As clearly demonstrated by Fritz et al. (2025), using lagged statistics that combine past and present edges can misrepresent the effects of SBT. For example, treating the presence of mutual friends or enemies at time t1t-1 as an exogenous covariate influencing the sign structure at time $t$ overlooks the fact that such covariates may not persist at time $t$, and may evolve in conjunction with other edge signs to form new balanced configurations. For this reason, in order to effectively model SBT effects, it is necessary to use endogenous network statistics that are implicitly dynamic, that is, they capture temporal dependencies without explicitly referencing past time points. It is important to emphasise that this reasoning also applies in our context, where SBT effects are estimated conditionally on the interaction structure. Indeed, in our case, assuming that such covariates remain constant would imply that both the interactions and their associated signs persist from time t1t-1 to time $t$, and are linked by an underlying structural dependency.
To test SBT effects, we adopt the geometrically-weighted statistics proposed by Fritz et al. (2025). These statistics can be included as binary terms $\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}}\right)$ and $\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}}\right)$ in 2-layer STERGMs (5) as the edgewise shared partner configurations shown in Figure 4 are used conditionally on the presence of interactions, resulting in binary configurations within a constrained support. For example, the geometrically weighted distribution of negative edgewise shared friends, $\texttt{gwesf}^{-}(\mathbf{y}_{t},\alpha)$, where $\alpha$ is the decay parameter, is equivalent to the binary geometrically weighted non-edgewise shared partners term, $\texttt{gwnsp}(\mathbf{z}_{t};\mathbf{x}_{t},\alpha)$, measured on $\mathbf{z}_{t}$ given $\mathbf{x}_{t}.$ In fact, conditional on $\mathbf{x}_{t}$, each dyad in $\mathbf{z}_{t}$ is binary, and therefore a negative edge that anchors the positive two-paths in the $\texttt{gwesf}^{-}$ configuration plays the same structural role as a non-edge anchoring the connected two-paths counted by the gwnsp configuration in standard binary networks.
@@ 8 Bayesian inference
We adopt a fully probabilistic Bayesian framework to estimate the 2-layer STERGM by defining a prior distribution over the model parameters $\pi(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}})$, such that the posterior distribution is given by:
@@
Algorithm 1 Approximate exchange algorithm for $\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})$
1:Initial values $\boldsymbol{\zeta}^{\mathcal{F}(0)},\boldsymbol{\zeta}^{\mathcal{P}(0)}$
2:Samples $\boldsymbol{\zeta}^{\mathcal{F}(i)},\boldsymbol{\zeta}^{\mathcal{P}(i)}$ for $i=1,\dots,I$ iterations
3:Compute $\mathbf{x}_{t}^{\mathcal{F}}$ and $\mathbf{x}_{t}^{\mathcal{P}}$ for all $t=1,\dots,T$
4:for $i=1$ to II do
5:  Propose $\boldsymbol{\zeta}^{\mathcal{F}^{\prime}}\sim h_{\mathcal{F}}(\cdot),$ and $\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\sim h_{\mathcal{P}}(\cdot)$
6:  for $t=1$ to $T$ do
7:   Simulate $\mathbf{z}_{t}^{\mathcal{F}^{\prime}}\sim p\left(\cdot\mid\mathbf{x}_{t}^{\mathcal{F}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}}\right)$
8:   Simulate $\mathbf{z}_{t}^{\mathcal{P}^{\prime}}\sim p\left(\cdot\mid\mathbf{x}_{t}^{\mathcal{P}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)$
9:   Set $\mathbf{z}_{t}^{\prime}=\mathbf{z}_{t}^{\mathcal{P}^{\prime}}\cup\left(\mathbf{z}_{t}^{\mathcal{F}^{\prime}}\backslash\mathbf{z}_{t}\right)$
10:  end for
11:  Compute acceptance probability:
12:
$\alpha=\min\left(1,\prod_{t=1}^{T}\frac{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\tilde{f}\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\times\frac{\pi\left(\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{\pi\left(\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\right)$ (7)
13:  With probability $\alpha$, set $(\boldsymbol{\zeta}^{\mathcal{F}(i)},\boldsymbol{\zeta}^{\mathcal{P}(i)})\leftarrow(\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}})$
14:  Otherwise, set $(\boldsymbol{\zeta}^{\mathcal{F}(i)},\boldsymbol{\zeta}^{\mathcal{P}(i)})\leftarrow(\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)})$
15:end for
To address the sampling challenges arising from highly correlated parameters, including correlations between the formation and persistence parameter groups, we employed an adaptive Metropolis–Hastings (AMH) proposal scheme (Haario et al., 2001). Although the adaptive direction sampling scheme (used by default in Caimo and Friel 2011) necessitates multiple chains and is less efficient over the long run, it remains valuable for sampling initial parameter values during the burn-in period.
@@
$\hskip 10.81218pt=\prod_{t=1}^{T}\frac{p\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}{p\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}\times\frac{p\left(\mathbf{z}^{\prime}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}(i-1)},\boldsymbol{\zeta}^{\mathcal{P}(i-1)}\right)}{p\left(\mathbf{z}^{\prime}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}^{\prime}},\boldsymbol{\zeta}^{\mathcal{P}^{\prime}}\right)}$
@@
A key advantage of this approach is that network sampling from the likelihood can be conducted with significantly fewer auxiliary iterations and lower computational cost, thanks to the sampling space constraint and reduced complexity of the statistics which are binary.